[figure: 9.7 cm]

MAGNETIC SUSCEPTIBILITY OF SINGLE CRYSTAL AND

POWDER TmFeO_{3}

Andrzej Bombik^{a} and Andrzej W. Pacyna^{b}

a Faculty of Phys. Nucl. Techniques, Mining and Metallurgy University, Kraków b Henryk Niewodniczański Institute of Nuclear Physics, Kraków, Poland

Low temperature and magnetic susceptibility of powder rare-earth orthoferrites depends essentially on technological history of compounds. Susceptibility in ZFC measurement differs from that received from FC one. Experiments show that temperature dependence of susceptibility difference χ_{FC} − χ_{ZFC} has the same singularities and overall temperature behaviour as the single crystal susceptibility measured along a direction. This coincidence suggests that phase transition temperatures in rare earth orthoferrites can be determined not only from single crystal investigation but also from data obtained for powder samples. This circumstance for example may be important in cases when proper single crystals are not available or when too weak magnetisation of materials does not permit to produce magnetic oriented powder samples. The careful study has been performed for TmFeO_{3} orthoferrites. The results of magnetic susceptibility investigation of powder samples (using Faraday method and DC magnetometer) are compared with the ones obtained for TmFeO_{3} single crystal. It is shown that phase transition temperatures in this orthoferrite determined from both methods coincide.

This work was supported by research grant No 2 P03B 035 14 from State Committee for Scientific Research.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Andrzej W. Pacyna

Address for correspondence : Instytut Fizyki Ja̧drowej im. Henryka Niewodniczańskiego ul. Walerego Eljasza Radzikowskiego 152 31-342 Kraków

Email address : Andrzej.Pacyna@ifj.edu.pl